OpenStudy (anonymous):

Given the function, f(x) = x2 – 3 find the value of f(3) am i supposed to plug the three into f(x) im so confused

OpenStudy (anonymous):

6 , 3 , 1 , -1 are my only answer options though.

OpenStudy (anonymous):

If there's answer options you could plug in the answers and see

OpenStudy (anonymous):

okay, so is it supposed to equal 3 when i plug a number into x2-3

OpenStudy (anonymous):

Wait, sorry, I said that wrong, you misunderstood, so did I, actually. f(x) = x^2 - 3 so f(3) = (3)^2 -3 = 9 - 3 = 6

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Yep :) Sorry, I confused us both at first